A piggy bank contains $4.40 in nickles, dimes and quarters. there are twice as many quarters as nickles and 5 more dimes than nickles. how many of each kind of coins are there?
Answer by ptaylor(2198)

Let's deal in pennies, ok
Let x=number of nickels; value of the nickels=5x
Then x+5=number of dimes; value of dimes=10(x+5)=10x+50
And 2x=number of quarters; value of quarters=25*2x=50x
Now we are told that the total value is 440, so:
5x+10x+50+50x=440 subtract 50 from each side
5x+10x+50x+50-50=440-50 collect like terms
65x=390 divide both sides by 65
x=6-------------------------number of nickels
x+5=6+5=11------------------------number of dimes
2x=2*6=12-------------------------number of quarters
CK
6*5+11*10+12*25=440
30+110+300=440
440=440

Hope this helps----ptaylor